OpenStudy (anonymous):

WILL MEDAL AND FAN!!!!!!!!!!!!!!! 9x^2-36

rishavraj (rishavraj):

a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

(3x+18)(3x-18)?

rishavraj (rishavraj):

18^2 = 36 ????

rishavraj (rishavraj):

\[\sqrt {36} = ????\]

OpenStudy (anonymous):

OHHH I divided by two my bad

rishavraj (rishavraj):

so wht it would be??

OpenStudy (anonymous):

it would be 6

OpenStudy (jeansweets):

324 for 18^2

rishavraj (rishavraj):

so whts the answer?

OpenStudy (jeansweets):

324

OpenStudy (anonymous):

(3x+6)(3x-6)

rishavraj (rishavraj):

@jeansweets 324 is the answer??? :/ hav a look over the question :)))) @Maddiem good work :)

OpenStudy (anonymous):

Thank you so much!!!!!!!

rishavraj (rishavraj):

u r most welcome :)))

OpenStudy (jeansweets):

use foil method maddiem 9x^2 +36 is for (3x+6)(3x-6)

OpenStudy (jeansweets):

if your asking for 18 to the power of 2 the answer is 324

OpenStudy (jeansweets):

@Maddiem I was answering @rishavraj question first